OpenStudy (anonymous):

find the area of the surface that is the graph of z=x+y^2 for 0<=x<=1,0<=y<=2

OpenStudy (anonymous):

\[\int\limits \int\limits f(x,y,g(x,y)) \sqrt((dg/dx)+(dg/gy)+1))\]

OpenStudy (anonymous):

i think this is the general equation that i need to use ... just dont know where or how to start off

OpenStudy (anonymous):

please help